OpenStudy (anonymous):

Simplify: (-2+3i)+(5+6i)

OpenStudy (anonymous):

3+9i

OpenStudy (anonymous):

Because -2 + 5 = 3, and 3i+6i=9i

mathslover (mathslover):

i will recommend u to see my tutorials of complex numbers

mathslover (mathslover):

\[\large{(-2+3i)+(5+6i)}\] \[\large{(-2+5)+i(3+6)}\] \[\large{3+9i}\]

mathslover (mathslover):

got it ?

OpenStudy (jiteshmeghwal9):

-2+3i+5+6i =3+9i U have to add only like terms.

OpenStudy (jiteshmeghwal9):

clear???????????